# Super-Kamiokande Strongly Constrains Leptophilic Dark Matter Capture in the Sun

### v1

@@ Abstract
To date, most terrestrial and celestial searches have focused on dark matter scattering with baryons, motivated by the similar masses of standard WIMPs and nuclei as well as the large enhancement in the interaction rate for spin-independent nuclear scattering. However, leptophilic dark matter models also provide well-motivated extensions of the standard model, which may be linked to several current anomalies in the neutrino and lepton sectors [14, 15, 16, 17, 18]. Leptophilic models for dark matter scattering motivate searches for the subsequent leptophilic annihilation of those same dark matter particles. Leptophilic annihilations can be highly advantageous for celestial body searches due to their high neutrino yield, potentially including direct annihilations to neutrino pairs.
The Sun is the preeminent astrophysical body for dark matter searches due to its proximity, which serves to both increase the neutrino flux observed at Earth, and also improve our background rejection, due to its motion across the sky [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. Solar neutrino observations have been used to probe novel physics, including constraining the dark matter/baryon cross-section across a wide range of dark matter masses.
In this letter, we use 10 years of Super-Kamiokande (Super-K) data to probe leptophilic dark matter capture and annihilation in the Sun. Figure 1 presents our main result, showing that current solar data rules out electron scattering cross-sections above $\sim$10-40 cm-2 for dark matter masses below 20 GeV across multiple leptonic annihilation channels. These constraints exceed the sensitivity of terrestrial experiments by more than an order of magnitude, and provide world-leading constraints on leptophilic dark matter scattering.
@@
Neutrinos from Dark Matter Annihilation — Dark matter capture by celestial bodies can produce regions inside the object with large dark matter densities. Dark matter in these regions can efficiently annihilate to standard model particles. In App. II we review the equations that govern annihilation inside a celestial body. However, when dark matter annihilation and capture is in equilibrium, the result simplifies to:
@@
where $L$ is the decay length of a mediator with energy $E_{\phi}\simeq m_{\chi}$. We note that this survival probability differs from the standard long-lived mediator case (e.g., [54]), which is calculated when the mediator decays to electromagnetic states. In that case, the electromagetic particles are absorbed by solar material if the mediator decays in the Sun, setting the survival probability to 0. Neutrinos, however, also usually escape from the Sun, removing this lower limit. Thus, long-lived mediators can produce a bright neutrino flux as long as they are produced on-shell, even if their decay length is short.
@@
Finally, while most neutrinos escape the solar interior, some can be absorbed by Hydrogen. In this work, we take this attenuation into account, which decreases the survival probability at higher neutrino energies. At $E_{\nu}=100$ GeV, this probability is $\simeq 96\%$, meaning that attenuation has a very small effect on our results through most of the relevant parameter space. We discuss the calculation for this probability in the Appendix III.
@@
Figure 1 (top) shows the full-sky atmospheric muon neutrino spectrum between 0.1–100 GeV (maroon), as calculated by the Super-K collaboration [66]. In order to determine the atmospheric neutrino background that could be confused with Solar neutrinos, we apply a cut based on the energy-dependent neutrino angular resolution of a water Cherenkov detector [67, 68, 69], which is a combination of two energy-dependent terms: an irreducible angular uncertainty correlating to the angular distribution of muons that are produced through muon neutrino scattering, and a detector uncertainty that depends on the ability of Super-K to reconstruct muon tracks. The combined angular resolution model is described in App. III.2, and is more accurate than previous studies [70], which assumed a constant angular resolution of  $20^{\circ}$, independent of neutrino energy.
@@
using the neutrino and water molecule cross section from Ref. [71], which we discuss in Appendix III.1. For the exposure of Super-K, we assume 22.5 kton of pure water with 10-year time scale. We also project Hyper-Kamiokande (Hyper-K) results using the same 10 yr timescale and a 187 kton mass.
@@
In Figure 4 we calculate the dark matter scattering limits for alternative dark matter/electron capture scenarios, including constant, velocity-dependent, and transfer momentum-dependent cross sections. We include the “Theory target” parameter spaces for sub-GeV leptophilic dark matter, with either a heavy  [75, 8, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87] or a light mediator  [75, 8, 88, 89, 90]. We note that terrestrial direct detection constraints on light mediators are much less sensitive, only reaching $\sim 10^{-35}$ cm2, which we show in Figure S6 of Appendix IV [91, 92, 93, 94, 95, 96].
@@
Conclusions — We use Super-K observations to strongly constrain leptophilic dark matter capture in the Sun. We study the neutrino signals from all leptophilic annihilation channels, finding that annihilations to $\tau^{+}\tau^{-}$, light-mediators, and the direct annihilation to neutrinos provide the strongest constraints. We obtain the strongest current limits on leptophilic dark matter scattering, reaching below 10-40 cm2 for dark matter masses between 4–20 GeV, and exceeding terrestrial direct detection searches by an order of magnitude between 4–100 GeV. We provide projections for future Hyper-K data, which can improve these constraints by an order of magnitude. By further extending our analysis down to a dark matter mass of 0.1 GeV and considering different dark matter capture scenarios, we demonstrate the potential of the Sun and other celestial bodies to probe the theory parameter space of dark matter/electron scattering [75, 8, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90]. Our results demonstrate the potential of neutrino detectors to investigate leptophilic dark matter, which strengthens the motivation for future experiments such as JUNO [98] and DUNE [99].
Acknowledgement — We thank Carlos Blanco, Chau Thien Nhan, and Ludwig Neste for fruitful discussions; Bei Zhou and Shirley Li for help with the background modeling of Super-K and Hyper-K; and Chris Cappiello for comments on the manuscript. TTQN thanks Jorge Martin Camalich and the Instituto de Astrofísica de Canarias for their hospitality. TTQN, TL and PC are supported by the Swedish Research Council under contract 2022-04283. TL is also supported by the Swedish National Space Agency under contract 2023-00242. This article/publication is based on the work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology). This work made use of Numpy [100], SciPy [101], matplotlib [102], Jupyter [103], as well as Webplotdigitizer [104].
@@ I.1 Constant cross section
$R^{-}_{\rm const}(w\to v)=\frac{2}{\sqrt{\pi}}\frac{\mu^{2}_{+}}{\mu}\frac{v}{w}n_{e}(r)\sigma_{\chi e}\Big{[}\chi(-\alpha,\alpha_{+})+\chi(-\beta_{-},\beta_{+})e^{\mu(w^{2}-v^{2})/u_{e}^{2}(r)}\Big{]}.$ (S9)
@@ I.5 Dark Matter Capture in the Sun and Jupiter
II Dark Matter Annihilation Rate Inside Celestial Bodies
@@
III Neutrinos interaction inside the Sun and Water Cherenkov Detectors
III.1 Neutrino cross section and attenuation
@@
III.2 Neutrino angular resolution and background cuts for Super-K
@@
IV Bounds on dark matter-electron cross section with Jupiter

### v2

@@ Abstract
To date, most terrestrial and celestial searches have focused on dark matter scattering with baryons, motivated by the similar masses of standard WIMPs and nuclei as well as the large enhancement in the interaction rate for spin-independent nuclear scattering. However, leptophilic dark matter models also provide well-motivated extensions of the standard model, which may be linked to several current anomalies in the neutrino and lepton sectors [14, 15, 16, 17, 18, 19, 20]. Leptophilic models for dark matter scattering motivate searches for the subsequent leptophilic annihilation of those same dark matter particles. Leptophilic annihilations can be highly advantageous for celestial body searches due to their high neutrino yield, potentially including direct annihilations to neutrino pairs.
The Sun is the preeminent astrophysical body for dark matter searches due to its proximity, which serves to both increase the neutrino flux observed at Earth, and also improve our background rejection, due to its motion across the sky [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Solar neutrino observations have been used to probe novel physics, including constraining the dark matter/baryon cross-section across a wide range of dark matter masses.
In this letter, we use 10 years of Super-Kamiokande (Super-K) observation to probe leptophilic dark matter capture and annihilation in the Sun. Figure 1 presents our main result, showing that current solar data rules out electron scattering cross-sections above $\sim$10-40 cm-2 for dark matter masses below 20 GeV across multiple leptonic annihilation channels. These constraints exceed the sensitivity of terrestrial experiments by more than an order of magnitude, and provide world-leading constraints on leptophilic dark matter scattering.
@@
We note that our capture rate exceeds that of Ref. [26] by a factor of $\sim$3 at low dark matter masses, and up to a factor of $\sim$7 at high dark matter masses, a fact first pointed out in Ref. [56], who obtain a smaller dark matter capture rate than in our work. In discussions with the authors of Ref. [56], we found that the difference between our results stems entirely from the choice in Ref. [56] to insert a kinematic cutoff into the formalism for the dark matter/electron scattering cross-section, which is valid only in the case that the electrons have zero temperature. If this cutoff is removed from the calculations of Ref. [56], they obtain results identical to ours. The initial kinematic constraint stems from Gould (1987) [57] who focused on nuclear scattering in the Earth, a scenario in which making a zero-temperature approximation is reasonable. However, applying such a cutoff (or similar variants of it) to high-temperature electrons artificially suppresses the dark matter capture rate by ignoring the extra stopping power of high-energy electrons that collide with the dark matter head on. We discuss this effect in further detail in the Supplemental Material.
Neutrinos from Dark Matter Annihilation — Dark matter capture by celestial bodies can produce regions inside the object with large dark matter densities. Dark matter in these regions can efficiently annihilate to standard model particles. In App. III we review the equations that govern annihilation inside a celestial body. However, when dark matter annihilation and capture is in equilibrium, the result simplifies to:
@@
where $L$ is the decay length of a mediator with energy $E_{\phi}\simeq m_{\chi}$. We note that this survival probability differs from the standard long-lived mediator case (e.g., [59]), which is calculated when the mediator decays to electromagnetic states. In that case, the electromagnetic particles are absorbed by solar material if the mediator decays in the Sun, setting the survival probability to 0. Neutrinos, however, also usually escape from the Sun, removing this lower limit. Thus, long-lived mediators can produce a bright neutrino flux as long as they are produced on-shell, even if their decay length is short.
@@
Finally, while most neutrinos escape the solar interior, some can be absorbed by Hydrogen. In this work, we take this attenuation into account, which decreases the survival probability at higher neutrino energies. At $E_{\nu}=100$ GeV, this probability is $\simeq 96\%$, meaning that attenuation has a very small effect on our results through most of the relevant parameter space. We discuss the calculation for this probability in the Appendix IV.
@@
Figure 1 (top) shows the full-sky atmospheric muon neutrino spectrum between 0.1–100 GeV (maroon), as calculated by the Super-K collaboration [71]. In order to determine the atmospheric neutrino background that could be confused with Solar neutrinos, we apply a cut based on the energy-dependent neutrino angular resolution of a water Cherenkov detector [72, 73, 74], which is a combination of two energy-dependent terms: an irreducible angular uncertainty correlating to the angular distribution of muons that are produced through muon neutrino scattering, and a detector uncertainty that depends on the ability of Super-K to reconstruct muon tracks. The combined angular resolution model is described in App. IV.2, and is more accurate than previous studies [75], which assumed a constant angular resolution of  $20^{\circ}$, independent of neutrino energy.
@@
using the neutrino and water molecule cross section from Ref. [76], which we discuss in Appendix IV.1. For the exposure of Super-K, we assume 22.5 kton of pure water with 10-year time scale. We also project Hyper-Kamiokande (Hyper-K) results using the same 10 yr timescale and a 187 kton mass.
@@
In Figure 4 we calculate the dark matter scattering limits for alternative dark matter/electron capture scenarios, including constant, velocity-dependent, and transfer momentum-dependent cross sections. We include the “Theory target” parameter spaces for sub-GeV leptophilic dark matter, with either a heavy  [78, 8, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92] or a light mediator  [78, 8, 93, 94, 95]. We note that terrestrial direct detection constraints on light mediators are much less sensitive, only reaching $\sim 10^{-35}$ cm2, which we show in Figure S7 of Appendix V [96, 97, 98, 99, 100, 101].
@@
Conclusions — We use Super-K observations to strongly constrain leptophilic dark matter capture in the Sun. We study the neutrino signals from all leptophilic annihilation channels, finding that annihilations to $\tau^{+}\tau^{-}$, light-mediators, and the direct annihilation to neutrinos provide the strongest constraints. We obtain world-leading limits on leptophilic dark matter scattering, reaching below 10-40 cm2 for dark matter masses between 4–20 GeV, and exceeding terrestrial direct detection searches by an order of magnitude between 4–100 GeV and project that Hyper-K data can improve these constraints by an order of magnitude.
By further extending our analysis down to a dark matter mass of 0.1 GeV and considering different dark matter capture scenarios, we demonstrate the potential of the Sun and other celestial bodies to probe the theory parameter space of dark matter/electron scattering [78, 8, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95]. Notably, our work obtains larger capture rates compared to previous work on the topic, which we find to be due to the extra stopping power of high-momentum electrons near the solar core. This result significantly enhances the expected sensitivity of all solar searches for dark matter electron scattering. Our results demonstrate the potential of neutrino detectors to investigate leptophilic dark matter, which strengthens the motivation for future experiments such as JUNO [103] and DUNE [104].
Acknowledgement — We thank Carlos Blanco, Chau Thien Nhan, and Ludwig Neste for fruitful discussions; Bei Zhou and Shirley Li for help with the background modeling of Super-K and Hyper-K; and Chris Cappiello, Dhashin Krishna, Ranjan Laha, Tarak Nath Maity, Nirmal Raj, Akash Kumar Saha, and Rinchen Sherpa for comments on the manuscript. We thank Raghuveer Garani, Joachim Kopp, and Sergio Palomares-Ruiz for helping us cross-check our results. TTQN thanks Jorge Martin Camalich and the Instituto de Astrofísica de Canarias for their hospitality. TTQN, TL and PC are supported by the Swedish Research Council under contract 2022-04283. TL is also supported by the Swedish National Space Agency under contract 2023-00242. This article/publication is based on the work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology). This work made use of Numpy [105], SciPy [106], matplotlib [107], Jupyter [108], as well as Webplotdigitizer [109].
@@ I.1 Constant cross section
$R^{-}_{\rm const}(w\to v)=\frac{2}{\sqrt{\pi}}\frac{\mu^{2}_{+}}{\mu}\frac{v}{w}n_{e}(r)\sigma_{\chi e}\Big{[}\chi(-\alpha_{-},\alpha_{+})+\chi(-\beta_{-},\beta_{+})e^{\mu(w^{2}-v^{2})/u_{e}^{2}(r)}\Big{]}.$ (S9)
@@ I.5 Dark Matter Capture in the Sun and Jupiter
II Comparison of Capture Rate Calculations with Previous Work –
On the Zero Temperature Approximation
After the initial submission of this manuscript, it was pointed out by Ref. [56] that the capture rates calculated in this work exceed those calculated by Ref. [26] by approximately a factor of three at low dark matter masses, and by a factor of up to seven for high dark matter masses. [^2] What follows in the remainder of this abstract is due to significant work in understanding the difference between this work and the initial version of the work in Ref. [56] (v1), Ref. [26] and Ref. [66]. We thank the authors of these manuscripts for their work in exchanging calculations and code in order to crystallize the key scientific issues which affected previous work on the topic. In this section, we explain the difference between these results, and show that our original results are correct because they account for the fact the Solar electron temperature has a sizable impact on the kinematics of the dark matter scattering cross-section. The results of our paper remain unchanged from the initial version of the manuscript. The enhancement of the dark matter capture rate demonstrated below also applies to Ref. [56].
The key difference between our results and Ref. [56] corresponds to an incorrect kinematic cutoff that has been added into the dark matter/electron scattering cross-section. This cutoff sets the total scattering rate to be zero in any regime where a zero-temperature electron would be unable to successfully capture the dark matter particle. In discussions with the authors of Ref. [56], we have confirmed that this is the primary difference between our results, and can confirm that our analyses reproduce each-other’s capture rates when this kinematic cutoff is added or removed. To further confirm that adding this term allows us to reproduce previous work, in Figure S4 we show a comparison between our capture rate and that of Ref. [26], showing that the arbitrary addition of this kinematic cutoff reduces our capture rates to identically match the solution in Ref. [26].
Figure S4: The total solar capture rate as a function of the dark matter mass and set at fixed cross sections of $\sigma_{\chi e}=10^{-40}$ cm2 and $\sigma_{\chi e}=10^{-42}$ cm2 from our current work (magenta lines), and that of Garani et al. [26] (red lines). Results are shown for constant cross-sections (left), velocity-dependent cross-sections (center) and momentum-dependent cross-sections (right). Our capture rate is larger by a factor of three for low dark matter masses and up to seven for high dark matter masses. Modifying our results by adding a kinematic cutoff that was based on a zero-temperature electron population (Equation S18) the green dotted points, which agrees with the result of Ref. [26], while adding an (unnecessary) kinematic cutoff corresponding to the modified high-temperature electron population (Equation S19) leads to the blue dotted points, and does not affect our result.
This zero-temperature kinematic cutoff was originally employed in the calculation of dark matter scattering with the Earth in Gould 1987 [57], in Equations 2.10 – 2.13. Gould also employs the zero-temperature approximation to simplify the expression from Equation 7.1 – 7.5, which is confusingly called the “finite temperature evaluation”, though it is no longer accurate in the case of finite-temperature electrons. Notably, however, this term does not appear in the simultaneously published Gould 1987 paper for dark matter scattering in the Sun [111].
While the zero-temperature approximation is very accurate for the 104 K temperatures in the Earth core, it produces sizable errors for the 107 K temperatures near the solar core, which correspond to the average electron moving at vth = $\sqrt{\frac{k_{b}T}{m_{e}}}\sim$ 0.1 c. The high-velocity of these electrons gives them extra stopping power in head-on collisions with more massive dark matter particles  [^3] It is also worth noting that Refs. [57, 111] were primarily concerned with nuclear scattering, where the effect of the Earth temperature on the scattering cross-section is even more negligible, justifying their zero-temperature approximation.
It is important to recognize the difference between the “average stopping power” of an electron and the “maximum stopping power” of an electron in a dark matter collision. The addition of a kinematic cutoff (given by a Heaviside function in the capture rate) is meant to represent the second case. In scenarios where a zero-temperature electron can slow down a dark matter particle below its escape velocity (for example at low dark matter masses), the total capture rate is maximized when all electrons are at zero temperature. Higher temperatures will decrease the dark matter capture rate by producing an electron sub-population that hits the dark matter particle from behind and further accelerates it.
However, in a scenario where a zero-temperature electron cannot slow down a dark matter particle to below its escape velocity (which is often the case for GeV-scale dark matter scattering off electrons in the Sun), the capture rate is largest when the electrons are at high temperature, and a sub-population of the electrons have the momentum necessary to slow down the dark matter particle in head-on collisions. Introducing a kinematic cutoff as soon as a zero-temperature electron cannot stop the dark matter particle ignores this large sub-population of electrons, which dominate the total capture rate for high-mass dark matter.
For clarity, we reproduce the calculation of the kinematic cutoff of Ref. [57] below. For a dark matter mass mχ and electron mass me, we first define the reduced masses:
$\mu=\frac{m_{\chi}}{m_{e}},$ (S14)
and:
$\mu_{\pm}=\frac{\mu\pm 1}{2}.$ (S15)
In a scenario where the electron has no momentum in the Sun’s reference frame, simple kinematics then tells us that the kinetic energy lost by the dark matter particle in the reference frame of the Sun is then uniformly distributed over the range:
$0\leq\frac{\Delta E}{E}\leq\frac{\mu}{\mu_{+}^{2}}.$ (S16)
Assuming the dark matter particle stars at velocity $w(r)$ where it is unbound, and must scatter to a velocity smaller than $v_{\rm esc}(r)$ where it is bound to the Sun, we must have:
$\frac{\Delta E}{E}\Big{|}_{\rm max}^{\rm zero-temp}\geq\frac{w^{2}-v_{esc}^{2}}{w^{2}}=\frac{u_{\chi}^{2}}{w^{2}},$ (S17)
by combining these two conditions we are left with a Kronecker delta function which can be applied to the scattering rate to ensure that the scattering results in a dark matter particle that is bound to the Sun.
$C_{\star}^{{\rm weak}}\propto\theta\left(\frac{\mu}{\mu_{+}^{2}}-\frac{u_{\chi}^{2}}{w^{2}}\right).$ (S18)
In the case where the electron has a high-momentum, this calculation clearly no longer applies, because it sets the capture rate to 0 based on a calculation that includes only the electron mass. If the electron has a significant momentum that is anti-parallel to the dark matter particle, conservation of momentum requires that it has additional stopping power capable of decreasing the dark matter velocity below vesc, leading to dark matter capture. For an electron with an initial velocity $-u_{e}$ (in the opposite direction of the dark matter), which collides with a dark matter particle with velocity $w$, one can calculate the maximum energy loss as:
$\frac{\Delta E}{E}\Big{|}_{\rm max}^{\rm finite-temp}=\frac{4m_{\chi}m_{e}}{(m_{\chi}+m_{e})^{2}}\frac{w+u_{e}}{w^{2}}\left(w-\frac{m_{e}}{m_{\chi}}u_{e}\right)=\frac{\mu}{\mu_{+}^{2}}\frac{w+u_{e}}{w^{2}}\left(w-\frac{m_{e}}{m_{\chi}}u_{e}\right).$ (S19)
Compared to the kinematic cutoff in Eq. S18, this maximum energy loss extends to significantly higher values in the case that $u_{e}$ is large, corresponding to an electron that is moving rapidly in the opposite direction of the dark matter particle.
However, we note that Equation S19 should also not be translated into a simple theta-function cutoff for the dark matter capture rate (although for the solar temperatures, dark matter masses, and dark matter velocities in our study, applying it artificially to our calculated capture rate has no effect on our results, as shown in Figure S4). There are two reasons why dark matter scattering off a high temperature-target cannot be limited by a single Kronecker-delta function cutoff: (1) there will be high-energy electrons up to arbitrarily high energies stemming from the tails of the Maxwell-Boltzmann distribution of the electron population, (2) the distribution of energy losses below the high-energy cutoff is no longer uniform, and is not represented by a Kronecker-delta function.
Fortunately, the original work of Gould (1987) [111] already correctly accounts for the full calculation of the capture rates for a Maxwellian distribution of dark matter particles scattering off of a Maxwellian distribution of standard model targets. The kinematics of the scatter, which enforces that the final velocity of the dark matter particle ($v$) falls below the escape velocity of the Sun is already correctly accounted for in Equation S2, by calculating the scattering probability $R_{e}^{-}(w\rightarrow v)$ and enforcing that the final value of $v$ falls below the escape velocity $v_{{\rm esc}}(r)$ by setting this as the maximum limit of integration in $dv$. Thus, we conclude that Equation S2, as discussed in the initial draft of this paper, provides the correct formalism for scattering off of high-temperature electrons in the Sun.
We note that applying our new results to several previous papers in the literature e.g.[66, 26, 27]) will enhance the capture rates in those calculations by factors of between 3–7, generally strengthening the conclusions of their work. Additionally, these results strengthen our projection for the sensitivity of future observatories (e.g., DUNE and HyperKamiokande) to dark matter/electron scattering.
III Dark Matter Annihilation Rate Inside Celestial Bodies
@@
IV Neutrinos interaction inside the Sun and Water Cherenkov Detectors
IV.1 Neutrino cross section and attenuation
@@
IV.2 Neutrino angular resolution and background cuts for Super-K
@@
V Bounds on dark matter-electron cross section with Jupiter